A plane can travel 470mph with the wind and 390mph against the wind. What is the speed of the plane if the air is still?
Found 3 solutions by josgarithmetic, lwsshak3, Alan3354:
Answer by josgarithmetic(39620)

Information is missing. What distances? Different or equal? What time quantities?

Give the complete exercise and someone could help.

Answer by lwsshak3(11628)

A plane can travel 470mph with the wind and 390mph against the wind. What is the speed of the plane if the air is still?
***
let x=speed of plane in still air
let c=speed of wind
x+c=speed of plane downwind
x-c=speed of plane against the wind
..
x+c=470
x-c=390
add
2x=860
x=430
What is the speed of the plane if the air is still? 430 mph

Answer by Alan3354(69443)

It's the average of the 2 groundspeeds, and it's called airspeed.